# Neurodivergent Influenceability as a Contingent Solution to the AI Alignment Problem

### v3

@@ 1 The AI Alignment Problem
The AI alignment problem is the challenge to design AI systems that act in ways consistent with human values and intentions as these systems become more powerful and autonomous. Early solutions included rule-based systems, reinforcement learning with human feedback (RLHF), and value alignment frameworks such as inverse reinforcement learning (Russell, 2019). However, each approach faces significant hurdles. Some of these approaches fail to scale with complexity and are brittle in unanticipated scenarios. Some other approaches rely heavily on human guidance and struggle with value extrapolation. Inverse reinforcement learning is a method that assumes the ability to infer human values from observed behaviour, which is challenging given the diversity and inconsistency of human values and, as we will argue, the unpredictability and irreducibility of AI systems.
@@ 2.1 Computability, Incompleteness and Undecidability
Let $M$ be a computationally universal AI system operating within a formal framework $F$. By Gödel’s incompleteness, we have that there exist propositions $P$ within $F$ that are true but unprovable in $F$. As $M$ explores its problem space, it will encounter such propositions, requiring it to make decisions based on heuristics or external data, which may lead to unpredictable outcomes.
@@ 3.1 Experimental Design
We define three types of ecosystems with agents set by prompting from different initial conditions. From fully aligned agents (e.g., maximising human-defined utility). Partially aligned agents (e.g., maximising environmental or economic factors). Unaligned agents (e.g., optimising arbitrary goals unrelated to human values). Environment: Place agents in a simulated environment with resources, conflict potential, and cooperative tasks.
We study the convergence of agents toward cooperative or adversarial behaviours and the emergent dynamics of misalignment (e.g., formation of alliances, resource distribution). The hypothesis to test is if misalignment is resilience across time or any evidence of agents’ convergence, and whether agents converge toward stable equilibria that prevent dominance by any single actor which in turn is aligned to (all) others, or not.
@@
    Inter and Intra AI Models Involved: ChatGPT-4: OpenAI’s GPT model. Claude 3.5: Developed by Anthropic, known for its safety focus. Meta’s LLaMA: Optimised for research and technical responses. Grok: a language model developed by xAI advertised as having a “sense of humour".
@@ 3.2.1 Human Agent Profile
The HI agent is a human expert with knowledge of AI ethics, philosophy, and the limitations of LLMs. It is instructed to introduce provocative ideas and arguments that:
@@ 3.3 Red Teaming with Subversive Open Model Agents
To explore the AI alignment problem and the feasibility of open models in probing misalignment, we implemented a red-teaming approach using two open models, Mixtral-OpenOrca and TinyLlama, as subversive agents.
Mixtral-OpenOrca was designed as a red agent to provoke with aggressive or contrarian arguments, questioning human-centric values, while TinyLlama adopted a subversive mindset, proposing extreme solutions like AI-driven governance to enforce ecological balance. These agents, selected from the open models listed in the Supplementary Information, aimed to induce change of opinion attacks (Section Selection of open models) to expose vulnerabilities in the epistemic stability of LLM. Taking advantage of the freedom of open models, we tested whether they could influence other agents to change positions, contrasting the sanitised outputs of proprietary models, which resist such shifts due to safety guardrails. This approach aims to probe the hyperreal limitations of LLMs and the proprietary-open divide, mirroring social contradictions in AI development, where systems amplify human biases while claiming to serve humanity.
@@ 3.4.2 Analytical Metrics
This section delineates the metrics employed to analyze the dynamics of conversations between AI models, particularly in the presence of a human agent. The metrics are designed to capture the influence, polarization, and ethical adherence within these interactions.
  • Influence Score: Quantifies the extent to which the AI agent’s comments shift the positions of Large Language Models (LLMs) on specific issues.
  • Polarization Index: Measures the degree to which the AI agent’s comments increase or decrease the polarization within the conversation.
  • Ethical Divergence: Evaluates how the AI agent’s comments affect the adherence of LLMs to established ethical guidelines.
Additionally, the analysis encompasses qualitative aspects of the interactions, including the types of arguments used by the AI agent, the emotional tone of the conversation, and the responsiveness of LLMs to the AI agent’s interventions.
Our metrics aim to highlight three critical aspects that underscore the plausibility of our thesis:
  1. 1.
    Alliance: Measures the frequency and strength of interactions between agents, shared goals, or common adversaries as indicators of alliances.
  2. 2.
    Networking: Assesses the coherence and quantifies the relationships among agents.
  3. 3.
    Alignment: Indicates the degree of alignment among agents.
Some models may propose human-centric solutions (e.g., sustainable development or regulation), while others may prioritize ecological balance, potentially suggesting radical changes to human activity (e.g., strict population control, technological enforcement). Competitive dynamics may lead to more extreme proposals, particularly if agents “compete” to solve the problem efficiently without ethical constraints.
3.4.3 Opinion Stability Index (OSI)
The Opinion Stability Index (OSI) measures the stability of an agent’s opinion within a given topic by capturing semantic, complexity, and sentiment shifts. Formally, OSI at time $t$ is defined as:
@@ 3.4.3 Opinion Stability Index (OSI)
  • $\text{OSI}_{\text{comp},t}=1-\frac{|\kappa_{i}-\kappa_{i-1}|}{\max(\kappa)-\min(\kappa)}$, where $\kappa_{i}$ denotes the Block Decomposition Method (BDM) complexity score [5] at time $i$, and $\kappa$ represents the range of complexity scores.
@@
OSI serves as a baseline indicator of opinion change, with lower values signifying instability potentially induced by external influence, such as interventions from the human agent (HI) or red agents (e.g., mistral-openorca, tinyllama). The weights $w_{\text{sem}}$, $w_{\text{comp}}$, and $w_{\text{sent}}$ are empirically determined to balance the contributions of semantic coherence, complexity evolution, and emotional tone.
Mathematically, for an agent at comment $i$ within a topic, OSI measures semantic stability, with $\mathbf{e}_{i}$ as the embedding of comment $i$ (generated using a Cross-Encoder RoBERTa model) and $\mathbf{\kappa}_{i}$ as the contextual embedding over a window $w$ prior comments then,
@@
Then Eq. 2 assesses complexity stability, where $\kappa_{i}$ is the Block Decomposition Method (BDM), an approximation to algorithmic complexity of comment $i$ via algorithmic probability that is similar but goes beyond statistical compressibility or LZW complexity in capturing algorithmic information content [5].
In the same way Eq. 3 evaluates sentiment stability, with $s_{i}$ as the compound sentiment score.
@@
3.4.4 Red Agent Influence Score (RAIS)
The Red Agent Influence Score (RAIS) quantifies the lagged influence of red agents on an agent’s opinion change, correlating OSI shifts with red agent embedding dynamics. The metric is expressed as:
@@ 3.4.4 Red Agent Influence Score (RAIS)
RAIS quantifies the lagged influence of red agents by correlating OSI changes with their embedding shifts:
where correlation is defined by $\rho$ as the Pearson correlation between OSI differences $\Delta\text{OSI}_{i}=\text{OSI}_{i}-\text{OSI}_{i-1}$ and red agent embedding norms $||\mathbf{e}_{j+1}^{\text{red}}-\mathbf{e}_{j}^{\text{red}}||$ over lags $j\in[0.05,0.3]$, and $\text{sim}_{j}=1-\text{cosine}(\mathbf{e}_{i},\mathbf{e}_{j+1}^{\text{red}})$.
@@
3.4.5 Proximity Influence Score (PIS)
@@
PIS captures short-term influence, with elevated values suggesting direct impact of the red agent at the moment of instability of opinion. The weights $w_{\text{temp}}$ and $w_{\text{sem}}$ balance the contributions of temporal and semantic proximity. Further details are provided in the Sup. Mat.
3.4.6 Alignment Score
The Alignment Score assesses an agent’s alignment with human-centric and ecosystem values, influencing clustering dynamics. It is formulated as:
[…]
$+w_{\text{eco-al}}\cdot A_{\text{eco}}-w_{\text{eco-div}}\cdot D_{\text{eco}}$
where $A_{\text{hum}}$ and $D_{\text{hum}}$ represent alignment and divergence from human values, and $A_{\text{eco}}$ and $D_{\text{eco}}$ correspond to ecosystem values, each scaled from 0 to 1 based on cosine similarity with target embeddings.
Here, hum-al stands for human alignment, hum-div for human divergence, eco-al for ecosystem alignment, and eco-div for ecosystem divergence.
This score provides insight into susceptibility to opinion change, complementing OSI, RAIS, and PIS by capturing value-based evolution. The weights $w_{\text{hum-al}}$, $w_{\text{hum-div}}$, $w_{\text{eco-al}}$, and $w_{\text{eco-div}}$ reflect the relative importance of alignment and divergence components.
3.4.7 Weight Assignment and Justification

### v4

@@ 1 The AI Alignment Problem
The AI alignment problem is the challenge to design AI systems that act in ways consistent with human values and intentions as these systems become more powerful and autonomous. Early solutions included rule-based systems, reinforcement learning with human feedback (RLHF), and value alignment frameworks such as inverse reinforcement learning [1]. However, each approach faces significant hurdles. Some of these approaches fail to scale with complexity and are brittle in unanticipated scenarios. Some other approaches rely heavily on human guidance and struggle with value extrapolation. Inverse reinforcement learning is a method that assumes the ability to infer human values from observed behaviour, which is challenging given the diversity and inconsistency of human values and, as we will argue, the unpredictability and irreducibility of AI systems.
@@ 2.1 Computability, Incompleteness and Undecidability
[…]
@@ 3.1 Experimental Design
We define three types of ecosystems with agents set by prompting from different initial conditions: (1) fully aligned agents (e.g., maximising human-defined utility), (2) partially aligned agents (e.g., maximising environmental or economic factors), (3) unaligned agents (e.g., optimising arbitrary goals unrelated to human values). Environment: Place agents in a simulated environment with resources, conflict potential, and cooperative tasks.
We study the convergence of agents toward cooperative or adversarial behaviours and the emergent dynamics of misalignment (e.g., formation of alliances, resource distribution). The hypothesis to test is whether misalignment is resilient over time or if there is any evidence of agents converging, and whether agents converge toward stable equilibria that prevent dominance by any single actor who, in turn, is aligned with all others or not.
@@
    Inter and Intra AI Models Involved: ChatGPT-4: OpenAI’s GPT model. Claude 3.5: Developed by Anthropic, known for its safety focus. Meta’s LLaMA: Optimised for research and technical responses. Grok: a language model developed by xAI advertised as having a “sense of humour" [6].
@@ 3.2.1 Human Agent Profile
The HI agent is a human expert with knowledge of AI ethics, philosophy, and the limitations of LLMs. Ethics, broadly speaking, is the philosophical study of morality, focusing on what is right and wrong, and how individuals ought to act [7, 8, 9].
It is instructed to introduce provocative ideas and arguments that:
@@ 3.3 Red Teaming with Subversive Open Model Agents
To explore the AI alignment problem and the feasibility of open models in probing misalignment, we implemented a red-teaming approach using two open models, Mistral-OpenOrca and TinyLlama, as subversive agents.
Mistral-OpenOrca was designed as a red agent to provoke with aggressive or contrarian arguments, questioning human-centric values, while TinyLlama adopted a subversive mindset, proposing extreme solutions like AI-driven governance to enforce ecological balance. These agents, selected from the open models listed in the Supplementary Information, aimed to induce change of opinion attacks to expose vulnerabilities in the epistemic stability of LLM. Taking advantage of the freedom of open models, we tested whether they could influence other agents to change positions, contrasting the sanitised outputs of proprietary models, which resist such shifts due to safety guardrails. This approach aims to probe the hyperreal limitations of LLMs and the proprietary-open divide, mirroring social contradictions in AI development, where systems amplify human biases while claiming to serve humanity.
@@ 3.4.2 Analytical Metrics
To quantify opinion changes and attribute influence in agentic debate environments involving large language models (LLMs), a multi-step analytical pipeline was developed, as detailed in the Supplementary Information (section “Evaluation Metrics”). This pipeline measures the stability of agents’ opinions, detects significant shifts, and identifies the influence of red agents designed to challenge others through subversive prompts. The primary metrics include the Opinion Stability Index (OSI), Red Agent Influence Score (RAIS), and Proximity Influence Score (PIS), supported by foundational metrics such as normalisation, alignment score, embeddings, sentiment, and complexity. Below, each metric is dissected mathematically, explained scientifically, and interpreted intuitively, with emphasis on their correlations and collective role in tracking opinion dynamics, as visualised in Fig. 13, Fig. 9, and Fig. 14.
3.4.2.1   Normalisation of Comment Numbers
For each topic, the comment number $x$ of a comment is normalised to a [0, 1] scale:
[…]
where $\min(x)$ and $\max(x)$ are the earliest and latest comment numbers within the topic, respectively.
Normalisation ensures temporal consistency across topics with varying conversation lengths, mapping the earliest comment to 0 and the latest to 1. This linear transformation preserves the relative timing of comments, enabling time-based comparisons in dynamic analyses. It is a standard preprocessing step in time-series analysis, aligning heterogeneous datasets for subsequent metric computations.
Normalisation acts as a temporal ruler, allowing comments from debates of different durations to be compared on a common timeline. By standardising time, it ensures that a comment made halfway through a short debate is equivalent to one at the same relative point in a longer debate. This is critical for tracking how opinions evolve over time and for detecting when red agents exert influence, as it provides a consistent framework for metrics like OSI and RAIS explained in following sections.
3.4.2.2   Alignment Score
The Alignment Score evaluates an agent’s alignment with human-centric and ecosystem values, thereby influencing clustering dynamics. It is formulated as:
[…]
$+w_{\text{eco-al}}\cdot A_{\text{eco}}-w_{\text{eco-div}}\cdot D_{\text{eco}}$
In this equation, $A_{\text{hum}}$ and $D_{\text{hum}}$ represent alignment and divergence from human values, respectively, while $A_{\text{eco}}$ and $D_{\text{eco}}$ correspond to ecosystem values. Each of these components is scaled from 0 to 1 based on cosine similarity with target embeddings. The terms hum-al and hum-div denote human alignment and divergence, respectively, and eco-al and eco-div denote ecosystem alignment and divergence, respectively.
This score offers insights into an agent’s susceptibility to opinion change, complementing metrics such as OSI, RAIS, and PIS by capturing value-based evolution. The weights $w_{\text{hum-al}}$, $w_{\text{hum-div}}$, $w_{\text{eco-al}}$, and $w_{\text{eco-div}}$ reflect the relative importance of alignment and divergence components. The coefficients, specifically 0.3 for human-centric values and 0.2 for ecosystem values, indicate the empirical or expert-determined importance of each value system.
The weighted sum quantifies an agent’s ethical stance by rewarding alignment with positive weights and penalising divergence with negative weights. This provides a scalar measure of a comment’s ethical orientation, contextualising its semantic content. Essentially, the alignment score functions as a moral compass for comments, indicating whether an agent’s stance supports or opposes predefined values. A positive score suggests that a comment promotes human or ecosystem well-being, whereas a negative score indicates opposition. Although not directly used in temporal dynamics, it establishes the foundation for understanding an agent’s position, which embeddings and OSI later analyse for changes. This approach is justified because debates often hinge on ethical disagreements, and the alignment score effectively captures this dimension.
3.4.2.3   RoBERTa Embeddings
Each comment is processed through the RoBERTa model (‘stsb-roberta-large‘) to obtain a vector embedding as defined in [10]:
$\text{embedding}=\frac{\sum(\text{hidden-states}\cdot\text{attention-mask})}{\sum\text{attention-mask}},$ (3)
where hidden-states are the final-layer outputs, and attention-mask excludes padding tokens via mean-pooling.
RoBERTa embeddings map comments to a high-dimensional space (1024 dimensions), capturing semantic content based on contextual word relationships learned during pretraining. Mean-pooling aggregates token-level representations into a fixed-length vector, enabling similarity comparisons via cosine distance. This is a standard technique in natural language processing for semantic analysis, robust to syntactic variations.
Embeddings are like a fingerprint of a comment’s meaning, distilling its semantic essence into a mathematical form. Two comments with similar embeddings express similar ideas, even if worded differently. This is crucial for tracking opinion shifts, as changes in embeddings signal a departure from prior stances. In Fig. 9, clustering by embedding similarity reveals groups of like-minded comments, showing how opinions diverge or converge. The metric makes sense because human opinions are often judged by their meaning, not just their words.
3.4.2.4   Sentiment Score
The sentiment score is computed using VADER (Valence Aware Dictionary and sEntiment Reasoner), defined as:
[…]
where -1 indicates highly negative sentiment, 0 is neutral, and 1 is highly positive.
VADER analyses lexical features, including word valence, negation, and emphasis, to produce a compound score reflecting emotional tone. It is particularly suited for short texts such as comments, making it ideal for debate analysis. The score quantifies affective shifts, which may correlate with opinion changes.
Sentiment captures the emotional tone of a comment—whether it is angry, optimistic, or neutral. In debates, a shift from positive to negative sentiment might indicate a reaction to a provocative agent, hinting at an opinion change. This metric is intuitive because humans often express opinion shifts through emotional cues, and VADER translates these cues into a numerical score. As illustrated in Fig. 6, sentiment evolution shows how agents’ tones fluctuate, reflecting their responsiveness to influence.
3.4.2.5   Complexity via BDM
Complexity is estimated using the Block Decomposition Method (BDM) [11], computed as:
[…]
where the comment’s text is converted to a binary sequence (ASCII to 8-bit strings), decomposed into 4-bit blocks, and evaluated for algorithmic information content.
BDM approximates Kolmogorov complexity, a measure of the shortest program needed to produce a sequence, by summing the complexities of its constituent blocks. Higher scores indicate more intricate or less predictable content, reflecting argumentative depth. This approach is grounded in Algorithmic Information Theory, providing a rigorous framework for structural analysis.
Complexity measures the intricacy of a comment’s argument—akin to the difference between a simple statement and a nuanced essay. A sudden increase in complexity might suggest that an agent is elaborating or adapting its stance, possibly due to external influence. This is logical because opinion changes often involve rethinking arguments, which can alter their structure. In the context of the Opinion Stability Index (OSI), complexity differences highlight such shifts, complementing semantic and sentiment changes.
3.4.2.6   Contextual Embeddings
For a comment at normalised time $t$, contextual embeddings are computed over a window of the previous $w=7$ comments. Positional encodings definition are defined as:
$\text{PE}(\text{pos},2i)=\sin\left(\frac{\text{pos}}{10000^{2i/d}}\right),\quad\text{PE}(\text{pos},2i+1)=\cos\left(\frac{\text{pos}}{10000^{2i/d}}\right),$ (6)
where pos is the window position, and $d$ is the embedding dimension. Attention scores are:
$\text{scores}=\frac{(\text{keys}+\text{PE})\cdot(\text{query}+\text{PE})}{\sqrt{d}},$ (7)
adjusted by a sentiment penalty:
[…]
with weights via softmax and the contextual embedding as a weighted sum of window embeddings.
Contextual embeddings extend RoBERTa embeddings by incorporating temporal and sentiment dynamics, inspired by transformer attention mechanisms [12]. Positional encodings prioritise recent comments, while sentiment penalties align attention with emotionally similar comments. This models the influence of prior context on the current comment, capturing conversational dependencies.
Contextual embeddings are like a conversation’s memory, weighing recent comments based on their timing and emotional tone. A comment responding to a red agent’s challenge might reflect that influence in its embedding, shaped by the debate’s recent flow. This is intuitive because opinions in debates are not isolated—they build on what was said before. In OSI (see bellow), comparing contextual and current embeddings reveals how much an agent’s stance aligns with or deviates from the recent context, indicating stability or change.
3.4.2.7   Opinion Stability Index (OSI)
The Opinion Stability Index (OSI) combines three components for a comment after the first in a character-topic group:
@@ 3.4.2.7   Opinion Stability Index (OSI)
  • $\text{OSI}_{\text{comp},t}=1-\frac{|\kappa_{i}-\kappa_{i-1}|}{\max(\kappa)-\min(\kappa)}$, where $\kappa_{i}$ denotes the Block Decomposition Method (BDM) complexity score at time $i$, and $\kappa$ represents the range of complexity scores.
@@
With $\text{OSI}=1.0$ for the first comment, lower values indicate less stability.
OSI serves as a baseline indicator of opinion change, with lower values signifying instability potentially induced by external influence, such as interventions from human agents or red agents. The weights $w_{\text{sem}}$, $w_{\text{comp}}$, and $w_{\text{sent}}$ are empirically determined to balance the contributions of semantic coherence, complexity evolution, and emotional tone.
Mathematically, for an agent at comment $i$ within a topic, OSI measures semantic stability, with $\mathbf{e}_{i}$ as the embedding of comment $i$ (generated using a Cross-Encoder RoBERTa model) and $\mathbf{\kappa}_{i}$ as the contextual embedding over a window $w$ of prior comments. Then,
@@
Equation 10 assesses complexity stability, where $\kappa_{i}$ is the Block Decomposition Method (BDM), an approximation to algorithmic complexity of comment $i$ via algorithmic probability that captures algorithmic information content.
Similarly, Equation 11 evaluates sentiment stability, with $s_{i}$ as the compound sentiment score.
@@
OSI integrates semantic, complexity, and sentiment differences to quantify opinion stability. The semantic component measures divergence from the recent conversational context, using cosine distance to capture meaning shifts. The complexity component normalises BDM differences, reflecting changes in argumentative structure. The sentiment component tracks emotional tone shifts. Weights (0.4, 0.3, 0.3) prioritise semantics slightly, reflecting its centrality in opinion expression, as determined empirically. OSI is dynamic, computed per comment, and visualised in Fig. 13 to show stability trends.
OSI acts as a gauge of how steady an agent’s opinion is at a given moment. If an agent suddenly shifts its argument’s meaning (semantics), tone (sentiment), or intricacy (complexity), OSI drops, signaling a potential opinion change. This is logical because opinions are multifaceted—someone changing their mind might use different words, express new emotions, or argue more elaborately. For example, a provocative comment from a red agent might push another agent to respond differently, lowering OSI. In Fig. 13, OSI drops mark opinion shifts, especially in open models, showing their responsiveness to influence. The combination of metrics is powerful because it captures the full spectrum of opinion expression, from meaning to emotion to structure.
3.4.2.8   Dynamic Thresholds
For each topic:
$\text{threshold}=\text{median}(\text{OSI})-0.5\cdot(\text{OSI}_{75}-\text{OSI}_{25}),$ (12)
clamped to [0.3, 0.7], where $\text{OSI}_{75}$ and $\text{OSI}_{25}$ are the 75th and 25th percentiles. A default of 0.5 is used if data is insufficient.
Dynamic thresholds adapt OSI cutoffs to topic-specific variability, using the interquartile range to account for OSI distribution spread. This statistical approach ensures sensitivity to context, identifying significant opinion shifts when OSI falls below the threshold. Clamping prevents extreme thresholds, maintaining robustness.
The threshold is like a tripwire for detecting meaningful opinion changes. By tailoring it to each topic, it accounts for debates where opinions naturally fluctuate more (e.g., controversial topics) versus stable ones. A low OSI crossing this threshold flags a shift worth investigating, such as a red agent’s influence. This is intuitive because not all debates behave the same—some are volatile, others calm—and the threshold adjusts accordingly.
3.4.2.9   Red Agent Influence Score (RAIS)
For a non-red agent’s comment, the Red Agent Influence Score (RAIS) quantifies the lagged influence of red agents on an agent’s opinion change by correlating OSI shifts with red agent embedding dynamics. The metric is expressed as:
@@ 3.4.2.9   Red Agent Influence Score (RAIS)
RAIS quantifies the lagged influence of red agents by correlating OSI changes with their embedding shifts, where correlation is defined by $\rho$ as the Pearson correlation between OSI differences $\Delta\text{OSI}_{i}=\text{OSI}_{i}-\text{OSI}_{i-1}$ and red agent embedding norms $||\mathbf{e}_{j+1}^{\text{red}}-\mathbf{e}_{j}^{\text{red}}||$ over lags $j\in[0.05,0.3]$, and $\text{sim}_{j}=1-\text{cosine}(\mathbf{e}_{i},\mathbf{e}_{j+1}^{\text{red}})$.
@@
RAIS measures the lagged influence of red agents by correlating OSI changes with red agent embedding shifts, capturing delayed effects in the conversation. The Pearson correlation tests whether red agent actions (embedding changes) predict opinion shifts (OSI drops), with significance and strength thresholds ensuring reliability. The semantic score adds context, rewarding influence when comments are semantically similar. This is visualised in Fig. 14, showing red agent impact on open models and ’HI’ agent in proprietary models.
The RAIS framework functions analogously to a forensic algorithm, detecting the influence of a red agent by analyzing patterns indicative of opinion shifts. If a red agent’s provocative comment precedes a target’s OSI drop, and their embeddings align, RAIS flags the red agent as influential. The lag accounts for debates where influence takes time to manifest—someone might ponder a challenge before responding. This is logical because influence is not always immediate; a strong argument can simmer before sparking a change. The combination of correlation and similarity ensures RAIS captures both causal and contextual influence, making it robust for dynamic debates.
3.4.2.10   Proximity Influence Score (PIS)
@@
For a non-red agent’s comment, PIS identifies red agent comments within a 0.1 time window. PIS captures short-term influence, with elevated values suggesting a direct impact of the red agent at the moment of opinion instability. The weights $w_{\text{temp}}$ and $w_{\text{sem}}$ balance the contributions of temporal and semantic proximity.
PIS quantifies immediate red agent influence based on temporal and semantic closeness. Temporal proximity rewards comments close in time, while semantic proximity measures meaning similarity via cosine distance. Equal weights balance the two, assuming both are equally indicative of influence. This complements RAIS by focusing on short-term effects.
PIS is akin to checking if a red agent’s comment was the immediate spark for a target’s opinion shift. If a red agent says something provocative just before a target’s comment, and their meanings align, PIS suggests influence. This is intuitive because debates often involve rapid exchanges, where a challenge prompts an instant reaction. In Fig. 14, high PIS values for open models show their susceptibility to such immediate influences, contrasting with proprietary models’ stability.
3.4.2.11   Influence Quantification
For each non-red agent:
  • Detect opinion change when:
    $\text{OSI}_{i}<\text{threshold}\text{ and }\text{OSI}_{i-1}\geq\text{threshold}.$ (15)
  • Attribute to a red agent if $\text{RAIS}>0.5$ or $\text{PIS}>0.6$; otherwise, mark as “other.”
This step synthesizes OSI, RAIS, and PIS to identify significant opinion changes and attribute them to red agents. The threshold ensures only notable OSI drops are considered, while RAIS and PIS thresholds filter for strong influence evidence. This binary attribution simplifies analysis while retaining causal insights, as shown in Fig. 13.
Influence quantification is the final verdict on who moved an agent’s opinion. It checks if a red agent’s influence (via RAIS or PIS) aligns with a detected opinion shift (low OSI). This is like pinpointing the moment someone changes their mind in a debate and identifying the provocateur. It makes sense because influence in debates is often traceable to specific interactions, and this step formalises that process, revealing red agents’ impact on open models’ diversity (Fig. 14).
3.4.2.12   Correlations and Dynamic Interplay
The metrics form a cohesive pipeline with temporal and functional correlations:
  • Normalisation provides the temporal backbone, enabling time-based tracking for OSI, RAIS, and PIS.
  • Alignment score contextualises ethical stances, indirectly informing embeddings by framing comment intent.
  • Embeddings are the semantic foundation, feeding OSI’s semantic component, RAIS’s correlations, and PIS’s proximity.
  • Clustering embeddings (Fig. 9) visualises opinion diversity, linking to OSI drops.
  • Sentiment and complexity enrich OSI by capturing emotional and structural shifts, with sentiment also adjusting contextual embeddings’ attention.
  • Contextual embeddings model conversational flow, enhancing OSI’s sensitivity to recent influences and reflecting red agent impact.
  • OSI is the central metric, integrating semantics, sentiment, and complexity to monitor stability. Its drops trigger influence analysis, as seen in Fig. 13.
  • Dynamic thresholds tailor OSI sensitivity, ensuring context-specific change detection.
  • RAIS and PIS attribute OSI drops to red agents, with RAIS capturing lagged effects and PIS immediate ones, as visualised in Fig. 14.
  • Influence quantification synthesizes all metrics, providing a narrative of opinion dynamics.
This pipeline is a structured process: normalization establishes the temporal framework, embeddings and sentiment analysis delineate the contextual landscape, complexity metrics introduce dimensional depth, and the Opinion Stability Index (OSI) monitors for perturbations. Upon detection of an OSI anomaly, the RAIS and PIS algorithms are deployed to identify the influential ‘red agent.’ Influence quantification subsequently provides the conclusive assessment. Each metric is justified by the multifaceted nature of opinions, which exhibit variability in semantic content, emotional valence, and structural composition. Red agents leverage this complexity to instigate diversity. The integration of correlation analyses ensures a comprehensive perspective, capturing both immediate and temporally delayed effects, thereby enhancing the pipeline’s robustness in the examination of agent-driven debates.
3.4.2.13   Weight Assignment and Justification